Project

General

Profile

NDOS Commissioning Meeting 2012/03/26

Attendance: Jon, Gavin, Athans, Peter, Gary, Jan, Susan, Jarek, Minerba, Leon

Agenda:

  1. Beam and Detector Update
  2. Cosmic Rate vs. Time in NDOS (Gavin)
  3. Warm vs. Cold APD Operations at NDOS (Peter)
  4. DataCheck Report (Jan)
  5. Installation Plans for the week
  6. DAQ Update and Plans for the week
  7. DCS Update and Plans for the week
  8. AOB

Beam and Detector Report

  • Peter: Detector synch problems Friday night, was very difficult to get a run going.
  • Jon: we did not take good quality data over the weekend, shifters let runs go that were obviously unsynched.

Cosmic Rate plots for blessing (Gavin)

  • Plot on slide 5 has been blessed!

Comparison of Cold and Warm Running for 5 APDs on NDOS (Peter)

  • Does cooling actually give us the desired effect? (lower noise, higher gain)
  • Gains of cooled APDs seems low. Leon: pretty much expected.
  • Jon: we should modify the voltage settings for the few APDs that are/will be cooled in the short-term. Will work on calculating correct settings automatically for the long-term.

DataCheck Report (Jan):

  • several runs
  • a hot FEB

Installation plans for this week

  • Leon: continuing purge system installation. Should finish this week, after which we'll be able to try to cool the newly installed APDs.

DAQ Update and Plans for next week

  • Peter: plan to roll out new software tomorrow, mostly changes to RunControl and ResourceManagement.
  • Jon: try partitioning later this week? Peter: perhaps.

DCS Update and Plans for next week

  • Athans: Nothing to add.

AOB

  • Nothing